OpenStudy (anonymous):

Can someone know if I've set up this problem right, I'm trying to find the derivative of this function using product rule (not chain rule ) h(t) = t squared ( 1 - t^2)

OpenStudy (anonymous):

So far I've gotten this: (1/2 t ^ -1/2) (1 - t ^ 2) + (t ^ 1/2) ( - t)

OpenStudy (amistre64):

what do you mean by "not chain rule"

OpenStudy (anonymous):

sorry that t ^ 2 should be square rooted

OpenStudy (anonymous):

Here is the derivative. h'(t) = 2t - 4t^3

OpenStudy (amistre64):

take a deep breathe, and do us all a favor; re write this so it resembles some sort of coherent mathematical notation :)

OpenStudy (anonymous):

alright hold on

OpenStudy (anonymous):

\[h(t) = \sqrt{t}(1-t ^{2})\]

OpenStudy (amistre64):

no matter how you do this; other than distributing the sqrt(t) thru the (..) you are going to have to implement the "chain rule" ....

OpenStudy (amistre64):

teh product rule is simply: rl = r'l + rl' ; where right is left and left is right :) just looks better to me like this ... r = sqrt(t) ; r' = 1/2(sqrt(t)) l = (1 - t^2) ; l' = -2t .. and thats where the chain rule would have been :)

OpenStudy (amistre64):

r = sqrt(t) ; r' = 1/2(sqrt(t)) l' = -2t l = (1 - t^2) ------------------------- \(-2t\sqrt{t}\) + \(\cfrac{(1-t^2)}{2\sqrt{t}}\)

OpenStudy (amistre64):

simplify as needed